OpenStudy (anonymous):

Evaluate the integral ∫(from 0 to 1) ∫(from y^2 to 1) ysinx^2 dxdy (Reverse the order of integration)

zepdrix (zepdrix):

\[\Large \int\limits_{y=0}^1\int\limits_{x=y^2}^1 y \sin (x^2)\;dx\;dy\] Mmm ok so to this, it'll help if we can draw our region of integration.There is our lower boundary on x, x=y^2.